OpenStudy (anonymous):

Which fraction is halfway between 3/5 and 7/10? 13/10 13/20 7/50 7/15

OpenStudy (anonymous):

You got an answer yet?

OpenStudy (triciaal):

express fractions with the same base to do comparison lcd of 5 and 10 is 10 3/5 = 6/10 and 7/10 = 7/10 halfway = (6+7)/2 divided by 10 = 6.5/10 = 65/100 reduced to the lowest term divide by 5 = 13/20 B

OpenStudy (anonymous):

Can I get the answer I am being timed?

OpenStudy (triciaal):

well you should stop long enough to read the answer

OpenStudy (anonymous):

I am sorry. Yes mam

OpenStudy (anonymous):

\[\frac{1}{2} \left(\frac{3}{5}+\frac{7}{10}\right)=\frac{13}{20} \]

OpenStudy (anonymous):

U go robtobey